# Diff4MMLiTS: Advanced Multimodal Liver Tumor Segmentation via Diffusion-Based Image Synthesis and Alignment

### v1

@@ Abstract
Index Terms:
Multimodal Learning, Liver Tumor Segmentation, Diffusion-Based Synthesis
I Introduction
@@
Liver tumor segmentation is crucial for clinical diagnosis and treatment. Previous studies have achieved notable results in unimodal liver tumor segmentation, confirming the effectiveness of deep learning techniques in this area [1],[2]. As different types of tumors exhibit distinct characteristics across various modalities, clinical radiologists often use multimodal CTs for more comprehensive diagnosis and more precise delineation of tumor boundaries [3]. Integrating multimodal information significantly enhances segmentation performance [4, 5]. However, compared to unimodal, the high annotation costs, privacy concerns, and data scarcity significantly constrain the exploration of multimodal segmentation methods.
@@
Recent multimodal learning studies have underscored the importance of data alignment [6, 7, 8]. Alignment is equally vital in multimodal medical image segmentation tasks. Public multimodal benchmarks like BraTS [9] and AutoPET [10] either employ standardized anatomical templates for rigid alignment or are collected simultaneously, inherently ensuring rigid data alignment. Thus, most current multimodal tumor segmentation methods operate under the assumption of strict data alignment, focusing primarily on multimodal fusion methods [11, 12]. In contrast, in real clinical settings, multimodal CT images often exhibit incomplete alignment or significant misalignment due to the spatiotemporal differences in data acquisition. Although some registration methods based on traditional paradigms (e.g., differential homogeneous deformation [13]) or deep learning [14, 15] have been proposed and achieved reasonable registration at the organ or tissue level, they still struggle to fully align the diffuse and indistinct tumor regions. This presents unavoidable noise for existing multimodal segmentation algorithms.
@@
Concurrently, due to the limited availability of images with tumors and the rising popularity of generative AI, recent studies have employed generative models (such as those based on Generative Adversarial Networks (GANs) or diffusion models to synthesize liver tumors [16, 17], demonstrating potential in enhancing tumor segmentation performance. However, these methods are confined to unimodal segmentation tasks and require normal CTs (without tumors) as synthesis conditions, somewhat restricting their application. Nevertheless, these synthetic tumor efforts provide two key insights: first, diverse tumor samples can be synthesized to alleviate the challenge of data scarcity and to enhance model generalization; second, synthesizing rigorously aligned multimodal CTs can partially mitigate the current dependence of segmentation algorithms on well-registered multimodal data.
@@
In such a context, we present an innovative and effective multimodal segmentation framework named Diff4MMLiTS. Specifically, the process begins with preliminary registration and data preprocessing, during which the target organ is roughly aligned, but the tumor remains misaligned. Therefore, we employ the mask of the annotated modality, dilating it to ensure adequate coverage of the tumor across various modalities. Subsequently, this mask is utilized in the Normal CT Generator (NCG) module for inpainting, yielding the normal CT. Within the latent space, the Multimodal CT Synthesizer (MCS) module leverages the multimodal CT features and randomly generated tumor masks to synthesize strictly aligned multimodal CTs with tumors. The aforementioned synthesis process employs the common mask, yet each modality’s CT is synthesized according to its unique feature distribution mapping. Finally, the roughly aligned real data is combined with the fully aligned synthetic data to train the segmentation model. Our contributions are as follows:
@@
Figure 1: The architecture of Diff4MMLiTS. Normal CT Generator (NCG) module uses the extended PVP mask to inpaint multimodal images to acquire normal CTs. Multimodal CT Synthesizer (MCS) module uses normal CTs to synthesize multimodal CTs. Multimodal Segmenter (MS) module trains segmenter using real and synthetic data. MCS comprises two components: a 3D autoencoder consisting of a CT feature encoder and decoder, a tumor synthesizer based on a diffusion model (DM).
@@
  • An innovative multimodal segmentation framework, Diff4MMLiTS, is proposed to be the first multimodal tumor segmentation framework based on diffusion synthesis as far as we know.
  • The Inpainting-Synthesis-Segmentation pipeline is designed to innovatively translate the challenges of tumor registration into a data synthesis problem, and the framework can be adapted to a variety of multimodal datasets without the need for perfect alignment.
  • Diff4MMLiTS is assessed on both internal and external datasets, with experimental results demonstrating that our method significantly surpasses existing multimodal segmentation techniques. Code and synthetic dataset will be available.
II Related Work
Multimodal segmentation aims to integrate information from different modalities through methods such as input fusion [2], feature interaction [12], and output fusion [4] to achieve better segmentation performance. For instance, mmformer[18] and RFNet[19] employ the transformer and CNN, respectively, to learn features across different modalities of brain tumors. However, common multimodal segmentation methods assume that the source images are strictly aligned at the pixel level. In contrast, in real clinical practice, multimodal images are often acquired at different times and using different imaging devices. Even with registration, perfect alignment is difficult to achieve, especially when dealing with localized lesions. To address this issue, researchers have developed novel methods for unregistered multimodal images, such as PAMRFuse[20] and Bsafusion[21]; however, these primarily focus on cross-modal fusion rather than multimodal segmentation. Our framework focuses on multimodal liver tumor segmentation, enhancing segmentation performance by addressing the alignment issue.
Medical image synthesis is also a highly regarded research topic, commonly applied in tasks such as missing modality completion, data augmentation, and data-efficient learning. Existing medical image synthesis methods are primarily based on GANs [22] and diffusion models [23],[24]. Some studies also explore tumor synthesis, such as Red-GAN[25] and LF-SynthSeg[26], focusing on synthesizing multimodal brain tumors. However, these methods are not specifically designed for liver tumor synthesis. Chen et al. [16] propose a diffusion-based method that successfully generates abdominal tumors. Wu et al. [17] introduce an adversarial training strategy that leverages unlabeled data for synthetic training. Although these methods are designed for liver tumors, they are constrained to unimodal synthesis and fail to exhibit the anticipated advantages when directly applied to multimodal tasks. To address these issues, we propose a synthetic image-based multimodal liver tumor segmentation framework that effectively tackles alignment issues in multimodal segmentation tasks by generating perfectly aligned multimodal data.
III Method
@@
Diff4MMLiTS comprises three pivotal components: the Normal CT Generator (NCG) module, the Multimodal CT Synthesizer (MCS) module, the Multimodal Segmenter (MS) module. These components are trained sequentially, with each preceding stage’s output serving as input for the next. The NCG module provides the CT background for the MCS module to synthesize CT images with tumor. The MCS module then supplies the MS module with synthetic images to train the segmenter. The following subsections will provide a detailed elaboration. The overall framework is shown in Fig. 1.
III-A Normal CT Generator (NCG) Module
@@
Since there is no publicly available paired multimodal normal CT dataset, the NCG module is used to generate paired multimodal normal CT data as the basis for tumor data synthesis. Inpainting involves reconstructing the missing or masked regions of an image [27]. Inspired by Suvorov [28], we employ an inpainting network based on Fast Fourier Convolution (FFC), which enhances the receptive field, ultimately achieving large mask inpainting with an image-wide receptive field. To reduce computational load, we perform slice-by-slice inpainting of 2D slices within 3D CT scans.
@@
Given an image $x\in\mathbb{R}^{H\times W\times 3}$ and a binary mask of the inpainting regions $m\in\mathbb{R}^{H\times W\times 1}$, stacking them will give: $x^{\prime}=\operatorname{stack}(x\odot m,m)$, where $x\odot m$ denotes the masked image. The feed-forward inpainting network $f_{\theta}(\cdot)$ then processes the input $x^{\prime}$ in a fully convolutional manner, then generates the inpainted image: $\hat{x}$ = $f_{\theta}(x^{\prime})$. Specifically, the image and mask pair serve as network inputs, which are downsampled into FFC residual blocks. During FFC processing, the input is divided into two parallel branches: a local branch, using traditional convolution, captures local information, while a global branch, based on real FFT, acquires global contextual information. These two outputs are then cross-fused and channel-wise concatenated to produce the final output.
@@
We fine-tune the model using pre-trained parameters and followed the training strategy of Suvorov et al [28]. For the fine-tuning of NCG, we employ original internal multimodal data to adapt it to the inpainting of multimodal CT images. After training, the inpainter is utilized to remove tumor foreground regions from multimodal CT with tumors. For each tumor region in the portal venous phase (PVP) mask, a 5×5 kernel is first employed for morphological closing, followed by a single iteration of dilation using a 3×3 kernel, to mitigate subjective noise and ensure maximal coverage of tumor regions across all modalities. NCG used the dilated PVP noise annotation as the mask to remove the foreground from labeled samples, ultimately producing a corresponding normal CT for each multimodal CT.
III-B Multimodal CT Synthesizer (MCS) Module
@@
Recently, diffusion models have achieved remarkable success in the synthesis of medical images. Inspired by Robin Rombach [29], we design a latent feature-based diffusion process capable of synthesizing multimodal liver tumors within latent space, significantly reducing computational complexity while maintaining synthesis quality. MCS is divided into two stages: first, a autoencoder is trained to learn comprehensive latent features, including the CT feature encoder $\mathcal{E}$ and decoder $\mathcal{D}$; second, a tumor synthesizer based on a latent diffusion model (LDM) is trained to generate synthetic images.
@@
Specifically, the VQGAN-based autoencoder performs the encoding and reconstruction between the image $x_{0}$ and its latent representation $z_{0}$, giving $\tilde{x}_{0}=\mathcal{D}(z_{0})=\mathcal{D}(\mathcal{E}(x_{0}))$. Since the universal autoencoding stage only requires training once and can be utilized repeatedly for multiple diffusion model trainings [29], we adapt the pre-trained autoencoder [16]. The diffusion process transpires within the latent space $Z$, where synthesized image features are reconstructed into the pixel space $X$.
The diffusion process comprises a forward process and a reverse process. The forward diffusion process transforms the data distribution into a standard Gaussian distribution at each time step $t\in\begin{bmatrix}1,\cdots,T\end{bmatrix}$ by the following process:
@@
III-C Multimodal Segmentation (MS) Module
The segmenter is a U-shape network. During training, MS dynamically synthesizes CTs with tumors from inpainted CTs, thereby constructing a hybrid dataset that combines synthetic and real data. The training objective of the segmenter is to minimize the voxel classification error relative to the ground truth. The loss function $\mathcal{L}_{seg}$ is consist of the Dice loss $\mathcal{L}_{dice}$ and the cross-entropy loss $\mathcal{L}_{ce}$ , which can be written as:
@@ III-C Multimodal Segmentation (MS) Module
where $\gamma$ is a factor for balancing the two loss contributions.
During inference, if the input does not consist of all four modalities, certain modalities will be duplicated to match the input size required by the model. For example, if only a single modality is present, it will be replicated four times and then concatenated before being fed into the segmentation model.
IV Experiments and Results
IV-A Datasets
To the best of our knowledge, there is currently no well-aligned publicly available multimodal liver tumor CT dataset. Therefore, we construct an in-house dataset, mmLiTs, based on multimodal abdominal CT data from Zhongda Hospital, Southeast University. After excluding cases with indistinct boundaries across all CTs, the final dataset includes 45 sets of four-phase abdominal CTs with in-plane resolutions of 512 × 512, spacing from 0.6445 mm to 0.9160 mm, slice counts ranging from 45 to 561, and thicknesses between 1.0 mm and 5.0 mm. Experienced radiologists manually annotated liver tumors, and the data are randomly split into training, validation, and test sets in a 3:1:1 ratio. Given that radiologists often use PVP images to identify tumor margins that are less discernible in other phases, the dataset provides annotations for PVP and attempts to register other modalities to PVP.
Additionally, we assess the generalizability of our method on the LiTS dataset [30]. Since the LiTS dataset is a unimodal dataset and does not involve multimodal alignment issues, we directly stack its unimodal images four times to match the input requirement of the segmentation model. We apply the same processing to all comparison methods, including Diff4MMLiTS, to ensure a fair performance comparison.
@@ IV-A Datasets
Table I: Quantitative Comparisons of Diff4MMLiTS with SOTA Methods on mmLiTs. The Best Results are Highlighted in Bold.
  Methods DSC(%) JAC(%) SE(%) PRE(%)
16’U-Net [1] 75.46 61.29 68.12 87.69
16’V-Net [31] 75.97 62.70 81.78 74.40
18’AttentionUNet [32] 67.80 53.75 65.25 77.72
18’nnUNet [2] 76.34 62.77 69.84 86.47
22’SwinUNETR [33] 77.28 63.41 71.72 86.02
18’SegResNet [34] 77.73 64.05 75.29 81.87
21’MAML [4] 76.06 63.08 67.92 91.27
22’mmformer [18] 77.32 64.32 78.76 77.94
23’A2FSeg [12] 69.74 55.90 61.93 88.66
24’PA-Net [5] 76.56 63.09 74.38 82.23
Diff4MMLiTS 79.02 66.33 74.47 85.24
IV-B Implementation Details
Diff4MMLiTS is implemented by PyTorch on a workstation equipped with NVIDIA RTX 3090. We use nnUNet as the default backbone. To avoid over-fitting, online data augmentations, including random flipping, rotation, and color jittering, are applied. All methods are trained for 1000 epochs. The preprocessing involves clipping the images to a [-21, 189] window and normalizing them to zero mean and unit standard deviation, followed by random cropping of 96×96×96 patches as input. The postprocessing filters the detection results by retaining only tumors located within the liver region. For the loss function, we set the value of $\gamma$ to 0.5.
The Dice Similarity Coefficient (DSC), Jaccard index (JAC), Sensitivity (SE), and Precision (PRE) are used to quantitatively evaluate the performance of Diff4MMLiTS and other state-of-the-art (SOTA) methods.
IV-C Overall Performance
To evaluate our method, we compare Diff4MMLiTS with various general and multimodal segmentation methods. Given that achieving precise liver segmentation is relatively straightforward, we focus specifically on the segmentation of intrahepatic tumors. The comparison results are presented in Table I, with visualizations shown in Fig. 2.
Table II: Comparison of Experimental Results on LiTS. The Best Results are Highlighted in Bold.

### v2

@@ Abstract
1 Introduction
@@
Liver tumor segmentation is crucial for clinical diagnosis and treatment. Previous studies have achieved notable results in unimodal liver tumor segmentation, confirming the effectiveness of deep learning techniques in this area [3dunet, nnunet]. As different types of tumors exhibit distinct characteristics across various modalities, clinical radiologists often use multimodal CTs for more comprehensive diagnosis and more precise delineation of tumor boundaries [zhang2023]. Integrating multimodal information significantly enhances segmentation performance [maml, mmformer, panet]. However, compared to unimodal, the high annotation costs and data scarcity significantly constrain the exploration of multimodal segmentation methods.
@@
Recent multimodal learning studies have underscored the importance of data alignment [prior, llava]. Alignment is equally vital in multimodal medical image segmentation tasks. Public multimodal benchmarks like BraTS [brats] and AutoPET [autopet] either employ standardized anatomical templates for rigid alignment or are collected simultaneously, inherently ensuring rigid data alignment. Thus, most current multimodal tumor segmentation methods operate under the assumption of strict data alignment [f2net, a2fseg]. In contrast, in real clinical settings, multimodal CTs often exhibit significant misalignment due to spatiotemporal differences in data acquisition. Although some registration methods based on traditional paradigms [fsl] or deep learning [cyclemorph, transmorph] have been proposed and achieved reasonable registration at the organ or tissue level, they still struggle to fully align the diffuse and indistinct tumor regions. This presents unavoidable noise for existing multimodal segmentation algorithms.
@@
In recent years, due to the improvement of generative models, tumor segmentation based on synthesis has become a highly regarded research topic [red, xpx]. Previous studies have demonstrated the effectiveness of using synthetic images to enhance liver tumor segmentation [syntumor, difftumor, freetumor]. However, these methods are primarily limited to unimodal images and fail to achieve expected performance when directly applied to multimodal tasks. Nevertheless, these efforts provide two key insights: first, synthetic tumor samples can be used to alleviate the challenge of data scarcity and enhance model generalization; second, synthesizing rigorously aligned multimodal CTs can partially mitigate the current dependence of segmentation algorithms on well-registered multimodal data.
@@
In this context, we present an innovative and effective multimodal segmentation framework named Diff4MMLiTS. Specifically, the process begins with preliminary registration and data preprocessing, during which the target organ is roughly aligned, but the tumor remains misaligned. Therefore, we employ the mask of the annotated modality, dilating it to ensure adequate coverage of the tumor across various modalities. Subsequently, this mask is utilized in the Normal CT Generator (NCG) module for inpainting, yielding the normal CT. Within the latent space, the Multimodal CT Synthesizer (MCS) module leverages the multimodal CT features and randomly generated tumor masks to synthesize strictly aligned multimodal CTs with tumors. The aforementioned synthesis process employs the common mask, yet each modality’s CT is synthesized according to its unique feature distribution mapping. Finally, the roughly aligned real data is combined with the fully aligned synthetic data to train the segmentation model. Our contributions are as follows:
@@
Figure 1: The architecture of Diff4MMLiTS. The NCG module uses the extended PVP mask to inpaint multimodal images to acquire normal CTs. The MCS module uses normal CTs to synthesize multimodal CTs. The MS module trains segmenter using real and synthetic data. MCS comprises two components: a 3D autoencoder consisting of a CT feature encoder and decoder, a tumor synthesizer based on a diffusion model.
@@
  • An innovative multimodal liver tumor segmentation framework, Diff4MMLiTS, is proposed to effectively leverage unaligned multimodal liver tumor images.
  • The designed Inpainting-Synthesis-Segmentation pipeline leverages a diffusion model to synthesize a large number of perfectly aligned multimodal liver tumor images, thereby significantly improving segmentation performance.
  • Diff4MMLiTS is evaluated on both internal and external datasets, with experimental results demonstrating outstanding performance.
2 Method
@@
Diff4MMLiTS comprises three pivotal components: the NCG module, the MCS module, and the MS module. These components are trained sequentially, with each preceding stage’s output serving as input for the next. NCG provides the CT background for MCS to synthesize CTs with tumor. MCS then supplies MS with synthetic images to train the segmenter. The following subsections will provide a detailed elaboration. The overall framework is shown in Fig. 1.
2.1 Normal CT Generator (NCG) Module
@@
Since there is no publicly available paired multimodal normal CT dataset, NCG is used to generate multimodal normal CT as the basis for tumor synthesis. Inspired by Suvorov [lama], we employ an inpainting network based on Fast Fourier Convolution (FFC). To reduce computational load, we perform slice-by-slice inpainting of 2D slices within 3D CT scans. Given an image $x$ and a binary mask of the inpainting regions $m$, they are stacked as follows:
$x^{\prime}=\operatorname{stack}(x\odot(1-m),m),$ (1)
@@
where $x\odot(1-m)$ denotes the masked image. The feed-forward inpainting network $f_{\theta}(\cdot)$ processes the input $x^{\prime}$ in a fully convolutional manner and generates the inpainted image $\hat{x}$ = $f_{\theta}(x^{\prime})$. Specifically, the image and mask pair serve as network inputs, which are downsampled into FFC residual blocks. During FFC processing, the input is divided into two parallel branches: a local branch, using traditional convolution, captures local information, while a global branch, based on real FFT, acquires global contextual information. These two outputs are then cross-fused and channel-wise concatenated to produce the final output.
@@
We fine-tune the model using pre-trained parameters and followed the training strategy of Suvorov et al. [lama]. For the fine-tuning of NCG, we employ original internal multimodal data to adapt it to the inpainting of multimodal CTs. After training, the inpainter is utilized to remove tumor foreground regions from multimodal CT with tumors. For each tumor region in the portal venous phase (PVP) mask, a 5×5 kernel is first employed for morphological closing, followed by a single iteration of dilation using a 3×3 kernel, to mitigate subjective noise and ensure maximal coverage of tumor regions across all modalities. NCG used the dilated PVP noise annotation as the mask to remove the foreground from labeled samples, ultimately producing a corresponding normal CT for each CT.
2.2 Multimodal CT Synthesizer (MCS) Module
@@
MCS is divided into two stages: first, a autoencoder is trained to learn comprehensive latent features, including the CT feature encoder $\mathcal{E}$ and decoder $\mathcal{D}$; second, a tumor synthesizer based on a latent diffusion model (LDM) [ldm] is trained to generate synthetic images. This latent feature-driven diffusion process enables efficient and reliable tumor generation within a compact latent space.
@@
Specifically, the VQGAN-based autoencoder performs the encoding and reconstruction between the image $x_{0}$ and its latent representation $z_{0}$, giving $\tilde{x}_{0}=\mathcal{D}(z_{0})=\mathcal{D}(\mathcal{E}(x_{0}))$. Since the universal autoencoding stage only requires training once and can be utilized repeatedly for multiple diffusion model trainings [ldm], we adapt the pre-trained autoencoder [difftumor]. The diffusion process comprises a forward process and a reverse process. The forward diffusion process transforms the data distribution into a standard Gaussian distribution at each time step $t\in\begin{bmatrix}1,\cdots,T\end{bmatrix}$ by the following process:
@@
2.3 Multimodal Segmentation (MS) Module
The segmenter is a U-shape network. During training, MS dynamically synthesizes CTs with tumors from inpainted CTs, thereby constructing a hybrid dataset that combines synthetic and real data. The training objective of the segmenter is to minimize the voxel classification error relative to the ground truth. The loss function $\mathcal{L}_{seg}$ is consist of the Dice loss $\mathcal{L}_{dice}$ and the cross-entropy loss $\mathcal{L}_{ce}$:
@@ 2.3 Multimodal Segmentation (MS) Module
where $\gamma=0.5$ is a factor for balancing the two loss contributions. During inference, if the input does not consist of all four modalities, certain modalities will be duplicated to match the input size required by the model.
3 Experiments and Results
3.1 Datasets
To the best of our knowledge, there is currently no publicly available well-aligned multimodal liver tumor CT dataset. Therefore, we construct an in-house dataset from the Second Affiliated Hospital of Southeast University, called mmLiTS. mmLiTS includes 45 sets of four-phase abdominal CTs with in-plane resolution of 512 × 512, spacing from 0.6445 mm to 0.9160 mm, slice counts ranging from 45 to 561, and thicknesses between 1.0 mm and 5.0 mm. Liver tumors are manually annotated by experienced radiologists, and the dataset is randomly split into training, validation, and test sets in a 3:1:1 ratio. Given that radiologists often rely on PVP images to identify tumor margins that are less discernible in other phases, annotations are provided for PVP. To further assess generalizability, we evaluate our method on the LiTS dataset [lits]. As LiTS is a unimodal dataset without alignment issues, we replicate each image four times to match the input format required by our segmentation model.
3.2 Implementation Details
Diff4MMLiTS is implemented by PyTorch on a workstation equipped with NVI-DIA RTX 3090. We use nnUNet as the default backbone. To avoid over-fitting, online data augmentations, including random flipping, rotation, and color jittering, are applied. All methods are trained for 1000 epochs. The preprocessing involves clipping the images to a [-21, 189] window and normalizing them to zero mean and unit standard deviation, followed by random cropping of 96×96×96 patches as input. The postprocessing filters the detection results by retaining only tumors located within the liver region.
Table 1: Quantitative comparisons of Diff4MMLiTS with state-of-the-art methods on mmLiTS. The best results are highlighted in bold.
  Methods DSC(%) JAC(%) SE(%) PRE(%)
16’U-Net [3dunet] 75.46 61.29 68.12 87.69
16’V-Net [vnet] 75.97 62.70 81.78 74.40
18’AttentionUNet [attentionnet] 67.80 53.75 65.25 77.72
18’nnUNet [nnunet] 76.34 62.77 69.84 86.47
22’SwinUNETR [swinunetr] 77.28 63.41 71.72 86.02
18’SegResNet [segresnet] 77.73 64.05 75.29 81.87
21’MAML [maml] 76.06 63.08 67.92 91.27
22’mmformer [mmformer] 77.32 64.32 78.76 77.94
23’A2FSeg [a2fseg] 69.74 55.90 61.93 88.66
24’PA-Net [panet] 76.56 63.09 74.38 82.23
Diff4MMLiTS 79.02 66.33 74.47 85.24
@@ 3.2 Implementation Details
3.3 Overall Performance
To evaluate our method, we compare Diff4MMLiTS with various general and multimodal segmentation methods. Given that achieving precise liver segmentation is relatively straightforward, we focus specifically on the segmentation of intrahepatic tumors. The comparison results are presented in Table 1, with visualizations shown in Fig. 2.
Compared to the U-Net and its variants, our method enhances the DSC by 1.74% over the second-best SwinUNETR. When compared with other multimodal medical image segmentation models, our method enhances DSC by 1.29% over the second best SegResNet. Moreover, Diff4MMLiTS significantly reduces false positives and false negatives, as reflected in the improved SE and PRE. This indicates enhanced tumor detection capability and a diminished likelihood of missed or incorrect detections. Overall, our proposed method has excellent superiority and achieves the best performance in the multimodal liver tumor segmentation task.
3.4 Generalizability and Adaptability of Diff4MMLiTS
We evaluate the proposed method on a publicly available external dataset to assess its out-of-distribution generalization capability, demonstrating that the model trained with Diff4MMLiTS performs effectively without retraining. All methods are trained on mmLiTS and tested on lesion samples from LiTS. The comparison results are shown in Table 2. Compared to nnUNet fully trained on real data, Diff4MMLiTS with the synthesis strategy achieves a 16.12% improvement in DSC. This implies that training models with such reliable synthetic images can effectively mitigate the risk of overfitting to in-distribution samples, thereby enhancing their ability to generalize to out-of-distribution samples.
To further assess the adaptability of Diff4MMLiTS, we integrate three different backbone models into the MS module, with the results summarized in Table 3. The results demonstrate that our framework adapts well to these backbones, yielding notable performance gains. Compared to models trained solely on real data, those using the hybrid training strategy achieve DSC improvements of 1.95%, 6.58%, and 2.68%, respectively. These findings highlight both the flexibility of our framework across different backbones and the effectiveness of the hybrid training strategy in enhancing segmentation performance.
Table 2: Comparison of experimental results on LiTS. The best results are highlighted in bold.